OpenStudy (anonymous):

integrate sin3theta*sin5theta dtheta. I've done the problem a few times but keep coming up with the wrong answer.

zepdrix (zepdrix):

What methods have you tried? Hmm I think `by parts` will work very nicely.

OpenStudy (anonymous):

I was using 1/2(cos(m-n)x-cos(m+n)x)

zepdrix (zepdrix):

Hmm I don't know what that is D:

OpenStudy (anonymous):

it's a set of three formulas we got in class. It was probably derived from integration by parts though originally

zepdrix (zepdrix):

grrr you should have rotated it :) lol

OpenStudy (anonymous):

haha, sorry

zepdrix (zepdrix):

Hmm that's a nifty little formula. So I guess that gives us,\[\large \int\limits \sin5\theta \sin3\theta\;d \theta \quad=\quad \frac{1}{2}\int\limits \cos2\theta-\cos8\theta\;d \theta\]right?

OpenStudy (anonymous):

I did a -2theta for m-n

zepdrix (zepdrix):

So you ended up with,\[\large \int\limits\limits \sin5\theta \sin3\theta\;d \theta \quad=\quad \frac{1}{2}\int\limits\limits \cos(-2\theta)-\cos8\theta\;d \theta\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Keep in mind that cosine is an `even function` so:\[\Large \cos(-2\theta) \quad=\quad \cos(2\theta)\]

OpenStudy (anonymous):

ah, okay, cool

zepdrix (zepdrix):

So what did you do from that step? Did you do a `u-substition` ? You'll want to try and get more comfortable with problems like this, so you don't have to bother with a u-sub. On integrals like this, we simply divide by the coefficient on theta when we integrate.

zepdrix (zepdrix):

Example:\[\Large \int\limits \cos(4\theta)\;d \theta \quad=\quad \frac{1}{4}\sin(4\theta)\] Understand what I mean? :o We just divide by the coefficient (the 4), instead of doing a fancy u-sub. If that's too confusing you can go ahead with a u-sub though :)

OpenStudy (anonymous):

that's what I was doing. The signs probably messed me up

OpenStudy (anonymous):

yeah, definitely. I don't mess with u subs on those simple ones

zepdrix (zepdrix):

\[\large \frac{1}{2}\int\limits\limits \cos2\theta-\cos8\theta\;d \theta \quad=\quad \frac{1}{2}\left(-\frac{1}{2}\sin(-2\theta)-\frac{1}{8}\sin8\theta\right)\]

zepdrix (zepdrix):

So you end up with something like that?

zepdrix (zepdrix):

Woops I meant to put cos(-2theta) in the original, to match what you've been working with. Same result though.

OpenStudy (anonymous):

that is what I've been getting

zepdrix (zepdrix):

So why do you think your answer is wrong? :3 maybe it just looks a little different than the book.

OpenStudy (anonymous):

:/

zepdrix (zepdrix):

From here, you would want to remember that sine is an `odd function`:\[\Large \sin(-2\theta) \quad=\quad -\sin(2\theta)\]

zepdrix (zepdrix):

So that's why that negative ends up disappearing on the first term.

OpenStudy (anonymous):

We do our homework online and it keeps counting my answer wrong. Online systems are so picky though. Glad I'm doing it right though

zepdrix (zepdrix):

Oh yah, that can be a bit tricky D:

zepdrix (zepdrix):

A bit frustrating i mean*

OpenStudy (anonymous):

Thank you for your help though. And btw, I've mastered that reduction formula now :P

zepdrix (zepdrix):

Oh cool c: